OpenStudy (anonymous):

Need help w/ some substitution... The integral (x+1)sqrt(2x+x^2)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}(x+1)\sqrt{2x+x^2}dx\]

OpenStudy (anonymous):

That might help...

OpenStudy (anonymous):

u=2x+x^2 du=2+2xdx du/2=1+xdx integral of 1/2sqrt(u)du=2/3(1/2)u^3/2 1/3(2x+x^2)^3/2+C

OpenStudy (anonymous):

I would set U = to 2x + \[x^{2}\]. You can then solve for x using that equation and plug it in where the (x+1) goes

OpenStudy (anonymous):

that didnt work out its supposed to be going down and easier to read sorry

OpenStudy (anonymous):

1/3(2x+x^2)^3/2+C

OpenStudy (anonymous):

Oh it's okay, I can figure out what you did :) Thanks!

OpenStudy (anonymous):

welcome